96m:20086 20J06

Siegel, Stephen F. ^{(}1-NW^{)}

The spectral sequence of a split extension and the cohomology of an extraspecial group of order p^{3} and exponent p. (English. English summary) J. Pure Appl. Algebra 106 (1996), no. 2, 185–198. This paper begins with a refinement of a theorem of Charlap and Vasquez describing the d_{2} differential in the spectral sequence for the cohomology of a split extention of finite groups, and a simplified proof of this theorem is given. The rest of the paper is devoted to the example where the group is extraspecial of order p^{3} and exponent p, and the subgroup is elementary abelian of order p^{3}. In this case, it turns out that E_{3} = E_{∞}. David Benson (1-GA)

\begin{matrix}c \\ ◯ \end{matrix} Copyright American Mathematical Society 1996